OpenStudy (anonymous):

Find the slope-intercept form of each equation;then graph each line. x + 3y = 6

OpenStudy (anonymous):

convert it to y=mx + b form

OpenStudy (anonymous):

i do not know how to do this can you please explain?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

so y=mx+b means where m equals the slope and b is the y-intercept

OpenStudy (anonymous):

so ur solving for "y" in order to put it in y=mx+b form

OpenStudy (anonymous):

therefore, you do:

OpenStudy (anonymous):



OpenStudy (anonymous):

so y=-1/3x+2